OpenStudy (anonymous):

x/(x+1)-x^2-1 = (-x^3-x^2+2x+1)/x+1 How do I get that answer (STEPS)

OpenStudy (madhu.mukherjee.946):

x/(x+1)-(x+1)(x-1)=(-x^3-x^2+2x+1)/x+1

OpenStudy (ryanwhite499):

Let's solve your equation step-by-step. x x+1 −x2−1= −x3−x2+2x+1 x +1 x x+1 +−x2+−1= −x3−x2+2x+1 x +1 Multiply all terms by (x+1)(x) and cancel: xx+(−x2)(x+1)(x)+−1(x+1)(x)=(−x3−x2+2x+1)(x+1)+1(x+1)(x) −x4−x3−x=−x4−2x3+2x2+4x+1(Simplify both sides of the equation) −x4−x3−x−(−x4−2x3+2x2+4x+1)=−x4−2x3+2x2+4x+1−(−x4−2x3+2x2+4x+1)(Subtract -x^4-2x^3+2x^2+4x+1 from both sides) x3−2x2−5x−1=0 (Use cubic formula) x=−1.285142,−0.221876,3.507019 Check answers. (Plug them in to make sure they work.) x=3.507019(Works in original equation) x=−1.285142(Works in original equation) x=−0.221876(Works in original equation) Answer: x=3.507019 or x=−1.285142 or x=−0.221876

OpenStudy (ryanwhite499):

@Dengeki

OpenStudy (anonymous):

@ryanwhite499 Thanks